OpenStudy (anonymous):

A line passes through the point (0, 2) and has a slope of - 1/4 . What is the equation of the line?

OpenStudy (shaik0124):

point slope form is y-y1=m(x-x1) y-2=-1/4(x-0) y-2=-1/4 x

OpenStudy (anonymous):

A. 4x +y = 2 B. x-4y=8 c. 4x-y= -2 d. x+4y =8

OpenStudy (shaik0124):

4y-8=-x x+4y=8

OpenStudy (shaik0124):

d is answer

OpenStudy (shaik0124):

got it?

OpenStudy (the_fizicx99):

@shaik0124 please refrain from giving out answer.

OpenStudy (shaik0124):

add +8+x on both sides to4y-8=-x

OpenStudy (shaik0124):

ook fizizzizizizziziziziziziziz haha :)